At 16 °C and 765 torr pressure, what will be the volume of 83 mol of NO gas? (Enter your answer in scientific notation.)
Answer #1

Given, temperature = T = 16℃ = (273+16)K = 289 K

Pressures = P =765 torr = 1 atm.

No of moles of NO gas = n = 83

Let volume of gas be V.

We know,

PV = nRT. (R is universal gas constant = .082 L atm/K mole)

=> V = nRT/P

=> V = 83 * .082 * 289/1

=> V = 1966.9 L = 1.967 * 10^3 L.